25. April 2016: Montag

LDAP/IdM:

  • Nightly Update: No new issues over the weekend.
  • Accounts:
    • No issues in 28 new accounts over the weekend.
    • VPN updates per request.
    • IAM removes one departmental account and renames a second to that same netid.
    • Group membership on zebra.
    • IAM removes one departmental account and routes its email to another one.
  • Development:
    • File that was updated on Friday had the wrong permissions and blew up the weekend.
    • Fixed that, and forced a full weekend update – and something is terribly wrong in the code.
      • Was not properly keeping track of when the look aside type changed. Fixed.

Backups:

  • Save Groups:
    • Two CoM/IS groups still running from the last Friday.
    • Failures: dpm1 (C disk), med23 (O disk)
  • Clones: Tape and Disk clones started.
  • Tapes:
    • Four (4) tapes reclaimed for the vaults.
    • Forty (40) tapes relabeled into the Scratch pool.
  • Issues:
    • EMC SR 78857308: ETA 481089: ETA is not complete – to be updated.

Annual Review

  • Did the Self-Review…

Tags:

Leave a Reply

You must be logged in to post a comment.